# Large Scale Diffusion Distillation via Score-Regularized Continuous-Time Consistency

### v1

@@ Abstract
This work represents the first effort to scale up continuous-time consistency distillation to general application-level image and video diffusion models. Although continuous-time consistency model (sCM) is theoretically principled and empirically powerful for accelerating academic-scale diffusion, its applicability to large-scale text-to-image and video tasks remains unclear due to infrastructure challenges in Jacobian–vector product (JVP) computation and the limitations of standard evaluation benchmarks. We first develop a parallelism-compatible FlashAttention-2 JVP kernel, enabling sCM training on models with over 10 billion parameters and high-dimensional video tasks. Our investigation reveals fundamental quality limitations of sCM in fine-detail generation, which we attribute to error accumulation and the “mode-covering” nature of its forward-divergence objective. To remedy this, we propose the score-regularized continuous-time consistency model (rCM), which incorporates score distillation as a long-skip regularizer. This integration complements sCM with the “mode-seeking” reverse divergence, effectively improving visual quality while maintaining high generation diversity. Validated on large-scale models (Cosmos-Predict2, Wan2.1) up to 14B parameters and 5-second videos, rCM matches or surpasses the state-of-the-art distillation method DMD2 on quality metrics while offering notable advantages in diversity, all without GAN tuning or extensive hyperparameter searches. The distilled models generate high-fidelity samples in only $1\sim 4$ steps, accelerating diffusion sampling by $15\times\sim 50\times$. These results position rCM as a practical and theoretically grounded framework for advancing large-scale diffusion distillation.
Figure 1: 5 random video samples from 4-step sCM, DMD2, and rCM on Wan2.1 1.3B. rCM resolves the quality issues of sCM while showing clear superiority to DMD2 in generation diversity.
@@ 1 Introduction
Our work represents the first effort to scale up continuous-time consistency distillation to general application-level image and video diffusion models. To this end, we design dedicated infrastructure by developing a FlashAttention-2 JVP kernel and enabling compatibility with parallelisms such as FSDP and CP. This allows us to explore sCM’s scaling behavior by applying it to 10B+ models and high-dimensional video data. Through this investigation, we reveal the quality issues of sCM in fine-detail generation and identify the error accumulation characteristic of CMs.
At a conceptual level, we argue that the nature of diffusion distillation methods is governed by their underlying divergence (Figure 2): forward (e.g., CMs), where the training examples are real or teacher-generated data, and reverse (e.g., score distillation), where the student is supervised only on self-generated samples. Forward divergence is known to encourage “mode-covering” by penalizing underestimation of any training sample likelihoods, which often results in spread-out densities and low-quality samples. In contrast, reverse divergence is inherently “mode-seeking” and beneficial to the visual quality of diffusion models (Zheng et al., 2025a), despite suffering in diversity. To remedy the quality issues of sCM, we propose integrating score distillation as a long-skip regularizer, which naturally complements sCM as they act on forward and reverse generation paths, respectively. We term the resulting framework, together with our other techniques like stable time-derivative computation, the score-regularized continuous-time consistency model (rCM).
rCM requires no engineering complexities such as multi-stage training, GAN tuning or extensive architecture/hyperparameter search. We validate its scalability on unprecedentedly large-scale models including Cosmos-Predict2 (NVIDIA, 2025) and Wan2.1 (Wan et al., 2025), covering T2I and T2V tasks up to 5 seconds and 14B parameters. Empirically, rCM matches or even surpasses DMD2 on quality metrics, while offering significant advantages in generation diversity. These results establish rCM as a promising and practical direction for large-scale diffusion distillation.
@@ 2.2 Consistency Models
Consistency models (CMs) (Song et al., 2023) aim to learn a consistency function $\bm{f}_{\theta}:(\bm{x}_{t},t)\mapsto\bm{x}_{0}$ which maps the point $\bm{x}_{t}$ at arbitrary time $t$ on the teacher PF-ODE trajectory to the initial point $\bm{x}_{0}$. The consistency function must satisfy the boundary condition $\bm{f}_{\theta}(\bm{x},0)\equiv\bm{x}$. To ensure unrestricted form and expressiveness of the neural network $\bm{F}_{\theta}(\bm{x}_{t},t)$, $\bm{f}_{\theta}$ is parameterized as $\bm{f}_{\theta}(\bm{x},t)=c_{\text{skip}}(t)\bm{x}+c_{\text{out}}(t)\bm{F}_{\theta}(c_{\text{in}}(t)\bm{x},c_{\text{noise}}(t))$ with $c_{\text{skip}}(0)=1,c_{\text{out}}(0)=0$. This parameterization aligns with practices in diffusion models (Karras et al., 2022). $\bm{f}_{\theta}$ is the direct counterpart of the clean data predictor (denoiser) in diffusion models, and typically initialized from the teacher denoiser $\bm{f}_{\text{teacher}}$. CM’s objective is to ensure consistent outputs at adjacent timesteps $t-\Delta t$ and $t$ on the teacher trajectory. Discrete-time CMs minimize the objective with $\Delta t>0$:
@@
Continuous-time CMs, represented by sCM (Lu & Song, 2024), offer a clean upgrade by taking the limit $\Delta t\rightarrow 0$. When $d(\bm{x},\bm{y})=\|\bm{x}-\bm{y}\|_{2}^{2}$, the CM loss simplifies to $\mathbb{E}_{\bm{x}_{0}\sim p_{\text{data}},\bm{\epsilon},t}\left[w(t)\bm{f}_{\theta}(\bm{x}_{t},t)^{\top}\frac{\mathrm{d}\bm{f}_{\theta^{-}}(\bm{x}_{t},t)}{\mathrm{d}t}\right]$, where $\frac{\mathrm{d}\bm{f}_{\theta^{-}}(\bm{x}_{t},t)}{\mathrm{d}t}=\nabla_{\bm{x}_{t}}\bm{f}_{\theta^{-}}(\bm{x}_{t},t)\frac{\mathrm{d}\bm{x}_{t}}{\mathrm{d}t}+\partial_{t}\bm{f}_{\theta^{-}}(\bm{x}_{t},t)$ is the tangent of $\bm{f}_{\theta}$ at $(\bm{x}_{t},t)$ along the teacher ODE trajectory $\frac{\mathrm{d}\bm{x}_{t}}{\mathrm{d}t}=\bm{v}_{\text{teacher}}(\bm{x}_{t},t)$. sCM employs the TrigFlow noise schedule $\alpha_{t}=\cos(t),\sigma_{t}=\sin(t)$ and preconditioning $c_{\text{skip}}(t)=\cos(t),c_{\text{out}}(t)=-\sin(t)$  [^1] There is a data std parameter $\sigma_{d}$ in original sCM formulation, inherited from EDM (Karras et al., 2022). For simplicity, we absorb it into $\bm{x}_{0}$ itself, i.e., define $\bm{x}_{0}\coloneqq\frac{\bm{x}_{0}^{\text{raw}}}{\sigma_{d}}$ for original data $\bm{x}_{0}^{\text{raw}}$., such that $\bm{F}_{\theta}$ is exactly the velocity predictor $\bm{v}_{\theta}$. The loss further reduces to [^2] For simplicity, we absorb $c_{\text{noise}}$ into $\bm{F}_{\theta}$ itself. $\mathbb{E}_{\bm{x}_{0}\sim p_{\text{data}},\bm{\epsilon},t}\left[\left\|\bm{F}_{\theta}(\bm{x}_{t},t)-\bm{F}_{\theta^{-}}(\bm{x}_{t},t)-w(t)\frac{\mathrm{d}\bm{f}_{\theta^{-}}(\bm{x}_{t},t)}{\mathrm{d}t}\right\|_{2}^{2}\right],$ where $\frac{\mathrm{d}\bm{f}_{\theta^{-}}(\bm{x}_{t},t)}{\mathrm{d}t}=-\cos(t)(\bm{F}_{\theta^{-}}(\bm{x}_{t},t)-\frac{\mathrm{d}\bm{x}_{t}}{\mathrm{d}t})-\sin(t)(\bm{x}_{t}+\frac{\mathrm{d}\bm{F}_{\theta^{-}}(\bm{x}_{t},t)}{\mathrm{d}t})$, and the full derivative $\frac{\mathrm{d}\bm{F}_{\theta^{-}}(\bm{x}_{t},t)}{\mathrm{d}t}=\nabla_{\bm{x}_{t}}\bm{F}_{\theta^{-}}(\bm{x}_{t},t)\frac{\mathrm{d}\bm{x}_{t}}{\mathrm{d}t}+\frac{\partial\bm{F}_{\theta^{-}}(\bm{x}_{t},t)}{\partial t}$ can be computed using forward-mode automatic differentiation, Jacobian-vector product (JVP). This objective is a simple MSE which enforces the instantaneous self-consistency at $(\bm{x}_{t},t)$. Recent works MeanFlow (Geng et al., 2025) and AYF (Sabour et al., 2025) are essentially a direct combination of sCM and consistency trajectory models (CTM) (Kim et al., 2023) under the rectified flow schedule (see Appendix E.1).
@@ 3.1 Algorithm Details
Adapting to Any Noise Schedule. sCM employs the TrigFlow forward process $\bm{x}_{t}=\cos(t)\bm{x}_{0}+\sigma_{t}\bm{\epsilon}$, while the teacher model is typically trained under other noise schedules such as rectified flow. Due to the equivalence between different noise schedules and parameterizations in diffusion models (Kingma et al., 2021; Zheng et al., 2023b), a TrigFlow-consistent wrapped teacher can be constructed without retraining. Specifically, let the teacher time be $t^{\text{raw}}$ with noise schedule $\alpha_{t^{\text{raw}}},\sigma_{t^{\text{raw}}}$. A reverse mapping $\phi$ (often analytic) from TrigFlow time $t$ to $t^{\text{raw}}$ can be derived by matching the signal-to-noise ratio, i.e., by solving $\frac{\sigma_{t^{\text{raw}}}}{\alpha_{t^{\text{raw}}}}=\tan(t)$. Denote $\bm{f}_{\text{teacher}}^{\text{raw}}(\bm{x}_{t^{\text{raw}}}^{\text{raw}},t^{\text{raw}})$ as the original teacher denoiser (can be transformed from other parameterizations). The wrapped teacher is
@@
Simplification. As our concerned models do not involve the unstable Fourier time embedding or AdaGN layers mentioned in sCM, but instead adopt positional time embedding, AdaLN, and QK normalization, we keep the network structure. Following sCM’s tangent normalization, the loss is
@@
where $p_{G}$ is a time distribution, $c=0.1$, and $\bm{g}=w(t)\frac{\mathrm{d}\bm{f}_{\theta^{-}}(\bm{x}_{t},t)}{\mathrm{d}t}$. Although BF16 avoids the overflow issues as in sCM’s FP16, we still follow the JVP rearrangement by setting $w(t)=\cos(t)$ and absorbing it into the JVP computation. sCM’s adaptive weighting, as also noted in AYF (Sabour et al., 2025), is unnecessary since $\left\|\bm{F}_{\theta}-\bm{F}_{\theta^{-}}-\frac{\bm{g}}{\|\bm{g}\|_{2}^{2}+c}\right\|_{2}^{2}=\frac{\|\bm{g}\|_{2}^{2}}{\|\bm{g}\|_{2}^{2}+c}\approx 1$ remains nearly constant.
@@ 3.3.2 Theoretical Analysis: Error Accumulation
$\frac{\mathrm{d}\bm{f}_{\theta^{-}}(\bm{x}_{t},t)}{\mathrm{d}t}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\cos(t)}(\bm{F}_{\theta^{-}}(\bm{x}_{t},t)-\bm{F}_{\text{teacher}}(\bm{x}_{t},t))-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sin(t)}(\bm{x}_{t}+\underbrace{\frac{\mathrm{d}\bm{F}_{\theta^{-}}(\bm{x}_{t},t)}{\mathrm{d}t}}_{\text{self-feedback (JVP)}})$ (5)

### v2

@@ Abstract
Although continuous-time consistency models (e.g., sCM, MeanFlow) are theoretically principled and empirically powerful for fast academic-scale diffusion, its applicability to large-scale text-to-image and video tasks remains unclear due to infrastructure challenges in Jacobian-vector product (JVP) computation and the limitations of evaluation benchmarks like FID. This work represents the first effort to scale up continuous-time consistency to general application-level image and video diffusion models, and to make JVP-based distillation effective at large scale. We first develop a parallelism-compatible FlashAttention-2 JVP kernel, enabling sCM training on models with over 10 billion parameters and high-dimensional video tasks. Our investigation reveals fundamental quality limitations of sCM in fine-detail generation, which we attribute to error accumulation and the “mode-covering” nature of its forward-divergence objective. To remedy this, we propose the score-regularized continuous-time consistency model (rCM), which incorporates score distillation as a long-skip regularizer. This integration complements sCM with the “mode-seeking” reverse divergence, effectively improving visual quality while maintaining high generation diversity. Validated on large-scale models (Cosmos-Predict2, Wan2.1) up to 14B parameters and 5-second videos, rCM generally matches the state-of-the-art distillation method DMD2 on quality metrics while mitigating mode collapse and offering notable advantages in diversity, all without GAN tuning or extensive hyperparameter searches. The distilled models generate high-fidelity samples in only $1\sim 4$ steps, accelerating diffusion sampling by $15\times\sim 50\times$. These results position rCM as a practical and theoretically grounded framework for advancing large-scale diffusion distillation.
Figure 1: 5 random video samples from 4-step sCM, DMD2, SiD, and rCM on Wan2.1 1.3B. rCM resolves the quality issues of sCM while showing clear superiority to DMD2/SiD in generation diversity, exhibiting highly similar object position/orientation/motion to teacher and sCM.
@@ 1 Introduction
Our work represents the first effort to scale up continuous-time consistency and JVP to general application-level image and video diffusion models. To this end, we design dedicated infrastructure by developing a FlashAttention-2 JVP kernel and enabling compatibility with parallelisms including FSDP and CP. This allows us to explore sCM’s scaling behavior by applying it to 10B+ models and high-dimensional video data. Through this investigation, we reveal the quality issues of sCM in fine-detail generation and identify the error accumulation characteristic of CMs.
At a conceptual level, we argue that the property of diffusion distillation methods is governed by their underlying divergence (Figure 2): forward (e.g., CMs), whose objectives are defined on real or teacher-generated data, and reverse (e.g., score distillation), where the student is supervised on self-generated samples. Forward divergence, commonly used in pre-training [^1] For example, MeanFlow aims to train a few-step model from scratch., is known to encourage “mode-covering” by penalizing underestimation of any training sample likelihoods, which often results in spread-out densities and low sample quality. In contrast, reverse divergence, commonly used in post-training, is inherently “mode-seeking” and excels in generation quality, despite suffering from mode collapse and low diversity.
Motivated by this complementarity, we address the quality limitations of sCM by integrating score distillation as a long-skip regularizer. This design naturally pairs with sCM: the two supervision signals operate on the forward (external) and reverse (self-generated) data paths, respectively. The broader philosophy of jointly leveraging forward and reverse divergences echoes recent work such as DDO (Zheng et al., 2025a), which incorporates reverse KL into maximum-likelihood training to achieve state-of-the-art FID on ImageNet, and DDRL (Ye et al., 2025), which incorporates the supervised fine-tuning stage into large-scale diffusion reinforcement learning. We term the resulting distillation framework, together with our other techniques like stable time-derivative computation, the score-regularized continuous-time consistency model (rCM).
rCM requires no engineering complexities such as multi-stage training, GAN tuning or extensive architecture/hyperparameter search. We validate its scalability on unprecedentedly large-scale models (Cosmos-Predict2 (Ali et al., 2025), Wan2.1 (Wan et al., 2025)), covering T2I and T2V tasks up to 5 seconds and 14B parameters. Empirically, rCM matches or even surpasses DMD2 on quality metrics, while mitigating mode collapse and offering notable advantages in generation diversity. These results establish rCM as a promising and practical direction for large-scale diffusion distillation.
Extension to Autoregressive Video Diffusion
The paradigm of rCM is also promising to autoregressive video diffusion (Yin et al., 2025) for interactive world models. In particular, the currently dominant approach, Self-Forcing (Huang et al., 2025), can be viewed as a well-instantiated reverse-KL-style DMD tailored to a bidirectional teacher and a causal student. rCM suggests that forward-divergence-based distillation by causal teacher with teacher forcing could potentially complement self-forcing and enhance diversity and motion dynamics, paving the way for future exploration.
@@ 2.2 Consistency Models
Consistency models (CMs) (Song et al., 2023) aim to learn a consistency function $\bm{f}_{\theta}:(\bm{x}_{t},t)\mapsto\bm{x}_{0}$ which maps the point $\bm{x}_{t}$ at arbitrary time $t$ on the teacher PF-ODE trajectory to the initial point $\bm{x}_{0}$. The consistency function must satisfy the boundary condition $\bm{f}_{\theta}(\bm{x},0)\equiv\bm{x}$. To ensure unrestricted form and expressiveness of the student neural network $\bm{F}_{\theta}(\bm{x}_{t},t)$, $\bm{f}_{\theta}$ is parameterized as $\bm{f}_{\theta}(\bm{x},t)=c_{\text{skip}}(t)\bm{x}+c_{\text{out}}(t)\bm{F}_{\theta}(c_{\text{in}}(t)\bm{x},c_{\text{noise}}(t))$ with $c_{\text{skip}}(0)=1,c_{\text{out}}(0)=0$. This parameterization aligns with practices in diffusion models (Karras et al., 2022). $\bm{f}_{\theta}$ is the direct counterpart of the clean data predictor (denoiser) in diffusion models, and typically initialized from the teacher denoiser $\bm{f}_{\text{teacher}}$. CM’s objective is to ensure consistent outputs at adjacent timesteps $t-\Delta t$ and $t$ on the teacher trajectory. Discrete-time CMs minimize the objective with $\Delta t>0$:
@@
Continuous-time CMs, represented by sCM (Lu & Song, 2024), offer a clean upgrade by taking the limit $\Delta t\rightarrow 0$. When $d(\bm{x},\bm{y})=\|\bm{x}-\bm{y}\|_{2}^{2}$, the CM loss simplifies to $\mathbb{E}_{\bm{x}_{0}\sim p_{\text{data}},\bm{\epsilon},t}\left[w(t)\bm{f}_{\theta}(\bm{x}_{t},t)^{\top}\frac{\mathrm{d}\bm{f}_{\theta^{-}}(\bm{x}_{t},t)}{\mathrm{d}t}\right]$, where $\frac{\mathrm{d}\bm{f}_{\theta^{-}}(\bm{x}_{t},t)}{\mathrm{d}t}=\nabla_{\bm{x}_{t}}\bm{f}_{\theta^{-}}(\bm{x}_{t},t)\frac{\mathrm{d}\bm{x}_{t}}{\mathrm{d}t}+\partial_{t}\bm{f}_{\theta^{-}}(\bm{x}_{t},t)$ is the tangent of $\bm{f}_{\theta}$ at $(\bm{x}_{t},t)$ along the teacher ODE trajectory $\frac{\mathrm{d}\bm{x}_{t}}{\mathrm{d}t}=\bm{v}_{\text{teacher}}(\bm{x}_{t},t)$. sCM employs the TrigFlow noise schedule $\alpha_{t}=\cos(t),\sigma_{t}=\sin(t)$ and preconditioning $c_{\text{skip}}(t)=\cos(t),c_{\text{out}}(t)=-\sin(t)$  [^2] There is a data std parameter $\sigma_{d}$ in original sCM formulation, inherited from EDM (Karras et al., 2022). For simplicity, we absorb it into $\bm{x}_{0}$ itself, i.e., define $\bm{x}_{0}\coloneqq\frac{\bm{x}_{0}^{\text{raw}}}{\sigma_{d}}$ for original data $\bm{x}_{0}^{\text{raw}}$., such that $\bm{F}_{\theta}$ is exactly the velocity predictor $\bm{v}_{\theta}$. The loss further reduces to [^3] For simplicity, we absorb $c_{\text{noise}}$ into $\bm{F}_{\theta}$ itself. $\mathbb{E}_{\bm{x}_{0}\sim p_{\text{data}},\bm{\epsilon},t}\left[\left\|\bm{F}_{\theta}(\bm{x}_{t},t)-\bm{F}_{\theta^{-}}(\bm{x}_{t},t)-w(t)\frac{\mathrm{d}\bm{f}_{\theta^{-}}(\bm{x}_{t},t)}{\mathrm{d}t}\right\|_{2}^{2}\right],$ where $\frac{\mathrm{d}\bm{f}_{\theta^{-}}(\bm{x}_{t},t)}{\mathrm{d}t}=-\cos(t)(\bm{F}_{\theta^{-}}(\bm{x}_{t},t)-\frac{\mathrm{d}\bm{x}_{t}}{\mathrm{d}t})-\sin(t)(\bm{x}_{t}+\frac{\mathrm{d}\bm{F}_{\theta^{-}}(\bm{x}_{t},t)}{\mathrm{d}t})$, and the full derivative $\frac{\mathrm{d}\bm{F}_{\theta^{-}}(\bm{x}_{t},t)}{\mathrm{d}t}=\nabla_{\bm{x}_{t}}\bm{F}_{\theta^{-}}(\bm{x}_{t},t)\frac{\mathrm{d}\bm{x}_{t}}{\mathrm{d}t}+\frac{\partial\bm{F}_{\theta^{-}}(\bm{x}_{t},t)}{\partial t}$ can be computed using the forward-mode automatic differentiation, Jacobian-vector product (JVP). This objective is a simple MSE which enforces the instantaneous self-consistency at $(\bm{x}_{t},t)$. Recent works MeanFlow (Geng et al., 2025) and AYF (Sabour et al., 2025) are essentially a combination of sCM and consistency trajectory models (CTM) (Kim et al., 2023) under the rectified flow schedule (see Appendix E.1).
@@ 3.1 Algorithm Details
Adapting to Any Noise Schedule. sCM employs the TrigFlow noise schedule $\bm{x}_{t}=\cos(t)\bm{x}_{0}+\sin(t)\bm{\epsilon}$, while the teacher model is typically trained under other schedules such as rectified flow. Due to the equivalence between different noise schedules and parameterizations in diffusion models (Kingma et al., 2021; Zheng et al., 2023b), a TrigFlow-consistent wrapped teacher can be constructed without retraining. Specifically, let the teacher time be $t^{\text{raw}}$ with noise schedule $\alpha_{t^{\text{raw}}},\sigma_{t^{\text{raw}}}$. A reverse mapping $\phi$ (often analytic) from TrigFlow time $t$ to $t^{\text{raw}}$ can be derived by matching the signal-to-noise ratio, i.e., by solving $\frac{\sigma_{t^{\text{raw}}}}{\alpha_{t^{\text{raw}}}}=\tan(t)$. Denote $\bm{f}_{\text{teacher}}^{\text{raw}}(\bm{x}_{t^{\text{raw}}}^{\text{raw}},t^{\text{raw}})$ as the original teacher denoiser (can be transformed from other parameterizations). The wrapped teacher is
@@
Simplification. As our concerned models do not involve the unstable Fourier time embedding or AdaGN layers mentioned in sCM, but instead adopt positional time embedding, AdaLN, and QK normalization, we keep the network structure. Following sCM’s tangent normalization [^4] MeanFlow’s adaptive weighting $w=1/(\|\Delta\|_{2}^{2}+c)^{p}$ under its best-performing $p=1$ is the same as tangent normalization., the loss is
@@
where $p_{G}$ is a time distribution, $c=0.1$, and $\bm{g}=w(t)\frac{\mathrm{d}\bm{f}_{\theta^{-}}(\bm{x}_{t},t)}{\mathrm{d}t}$. Although BF16 avoids the overflow issues as in sCM’s FP16, we still follow the JVP rearrangement by setting $w(t)=\cos(t)$ and absorbing it into the JVP computation [^5] We find that the weighting $\cos(t)$ and JVP rearrangement are also helpful under BF16.. sCM’s adaptive weighting, as also noted in AYF (Sabour et al., 2025), is unnecessary since $\left\|\bm{F}_{\theta}-\bm{F}_{\theta^{-}}-\frac{\bm{g}}{\|\bm{g}\|_{2}^{2}+c}\right\|_{2}^{2}=\frac{\|\bm{g}\|_{2}^{2}}{\|\bm{g}\|_{2}^{2}+c}\approx 1$ remains nearly constant.
@@ 3.3.2 Theoretical Analysis: Error Accumulation
$\frac{\mathrm{d}\bm{f}_{\theta^{-}}(\bm{x}_{t},t)}{\mathrm{d}t}=-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\cos(t)}(\bm{F}_{\theta^{-}}(\bm{x}_{t},t)-\bm{F}_{\text{teacher}}(\bm{x}_{t},t))-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sin(t)}(\bm{x}_{t}+\underbrace{\frac{\mathrm{d}\bm{F}_{\theta^{-}}(\bm{x}_{t},t)}{\mathrm{d}t}}_{\text{self-feedback (JVP)}})$ (5)